# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v1

@@ 3.4 Boundary Consistency Regularized Multi-Task Learning (BCRM)
Figure 3: Overview of the BoundMatch architecture processing labeled samples, illustrating multi-task learning (MTL) of boundary detection alongside segmentation. The Boundary-Semantic Fusion (BSF) module integrates learned boundary cues from the boundary head into the segmentation head’s features, promoting sharper segmentation contours. Concurrently, the Spatial Gradient Fusion (SGF) module refines the boundary predictions by fusing them with the spatial gradient of the segmentation mask ($\nabla M$). This refinement improves the quality of boundary supervision for the labeled loss and is crucial for generating cleaner boundary pseudo-labels used in the consistency regularization process for unlabeled data.
@@ 3.4.1 Notation and Boundary Detection Head
We adopt a lightweight boundary detection head inspired by SBCB [11] and CASENet [26]. The architecture leverages hierarchical features from the backbone (e.g., features from Stem, layer1, layer2, and layer4 of a ResNet). Features from the earlier layers are processed to yield single-channel outputs, while the layer4 feature produces an output with $C$ channels, where $C$ is the number of classes. Specifically, the features from these layers are passed through a single $3\times 3$ convolutional layer, respectively, each followed by BN and ReLU activation. These processed features are then bilinearly upsampled to the original image size. Following [11], we use an additional $3\times 3$ convolutional layer to reduce artifacts from upsampling. The outputs are then fused using a sliced concatenation and undergo a grouped convolution to form the final boundary prediction map of shape $C\times H\times W$.
Let $Q$ denote the set of boundary predictions (e.g., fused output $q_{\text{fuse}}$ and the layer4 output $q_{\text{last}}$). For the binary boundary detection task, the ground-truth boundary is denoted by $z$ and the boundary prediction $q$ is a sigmoid-activated probability map in $[0,1]$. We supervise these outputs using a reweighted pixel-wise binary cross-entropy loss:
@@
$\mathcal{L}_{bdry}^{L}=\sum_{q\in Q^{L}}\frac{1}{B^{L}}\sum_{b=1}^{B^{L}}\frac{1}{CHW}\sum_{c,i,j}H_{bdry}\Big{(}z(b,c,i,j),\,q(b,c,i,j)\Big{)}.$ (6)
@@ 3.4.2 MTL for SS-SS with Boundaries
For the labeled set, the boundary detection head is trained alongside the segmentation head. For unlabeled data, we perform consistency regularization analogous to segmentation as shown in Fig. 1. The teacher’s fused boundary is used to obtain a hard pseudo-label with threshold $\tau_{bdry}$: $\hat{q}^{T}=\mathbf{1}\Big{(}q^{T,w}\geq\tau_{bdry}\Big{)}$. The unsupervised boundary loss is then:
$\mathcal{L}_{bdry}^{U}\\ =\sum_{q\in Q^{U}}\frac{1}{B^{U}}\sum_{b=1}^{B^{U}}\frac{1}{CHW}\sum_{c,i,j}H_{bdry}\Big{(}\hat{q}^{T}(b,c,i,j),\,q^{S,s}(b,c,i,j)\Big{)}.$ (7)
@@ 3.5.1 Boundary-Semantic Fusion (BSF)
Inspired by GSCNN [10], BSF explicitly integrates boundary predictions into the segmentation head. The output from the boundary detection head is concatenated with the features from the Atrous Spatial Pyramid Pooling (ASPP) using the bottleneck layer of the DeepLabV3+ decoder. We emphasize that there are no extra layers are introduced by BSF, and we only add extra channels for the bottleneck layer. This explicit incorporation of boundary cues helps localize object contours more accurately in the final segmentation.
@@ 3.5.2 Spatial Gradient Fusion (SGF)
Figure 4: Visualization of the output of spatial gradient operator on segmentation prediction. The produced mask $\nabla M$ is clean, but instance insensitive boundaries.
@@
$[q_{\text{fuse},1},\,q_{\nabla M,1},\,q_{\text{fuse},2},\,q_{\nabla M,2},\,\dots,\,q_{\text{fuse},C},\,q_{\nabla M,C}],$ (9)
forming a $2C$-channel activation map. This map is then processed by a convolutional layer with $C$-grouped convolutions to yield a refined semantic boundary probability map $q_{\text{refine}}$. This grouped convolution refines the boundary predictions by considering the spatial gradients, effectively sharpening the boundaries between different semantic regions.
The refined boundaries are supervised using the same $H_{bdry}$ loss. Moreover, we introduce a duality loss on the labeled set to enforce consistency between the spatial gradient and the ground-truth boundaries:
@@
We apply the duality loss only on labeled samples because the ground-truth boundary annotations are reliable; applying it on unlabeled data, which relies on pseudo-labels, introduces instability.
@@ 3.6 BoundMatch
Although demonstrated primarily with our SAMTH baseline throughout the experiments, BoundMatch is designed to be flexible and compatible with various CR methods. For instance, BoundMatch can be integrated with UniMatch [9] by incorporating the boundary consistency loss alongside UniMatch’s existing feature-level and strong-view consistency losses. We observed that using only a single model (without a teacher) tends to generate noisy boundary pseudo-labels; therefore, employing an EMA-updated teacher model is crucial for stability and quality. Similarly, integration with PrevMatch [42] is achieved by computing the boundary consistency loss using its robust ensemble pseudo-labels, thereby circumventing potential instability associated with single-iteration predictions. Furthermore, the architectural simplicity of the BSF and SGF fusion modules facilitates the incorporation of BoundMatch’s boundary-enhancement components into diverse segmentation models, a versatility explored in our ablation studies.
@@ 4.1 Experimental Configuration
Here, we specify the experimental configuration used for evaluating BoundMatch. This includes the datasets utilized (covering urban scenes and standard SS-SS academic benchmarks), the model architectures, and key implementation details.
@@ 4.1.2 Implementation Details
Cityscapes, BDD100K, SYNTHIA: We use a base learning rate of $0.01$. Loss weights are set to $\lambda_{seg}=1.0$, $\lambda_{bdry}=0.5$, $\lambda_{dual}=1.0$, with an overall unsupervised loss weight $\lambda=1.0$. The EMA decay is $\alpha=0.99$, and the confidence threshold for segmentation pseudo-labels is $\tau=0$. Training uses a mini-batch size of 16 (8 labeled, 8 unlabeled) for 80,000 iterations. The input crop size is $801\times 801$ for Cityscapes and $641\times 641$ for BDD100K and SYNTHIA.
Pascal VOC: Hyperparameters are adjusted: base learning rate is $0.001$, boundary loss weight $\lambda_{bdry}=0.1$, duality loss weight $\lambda_{dual}=0.5$ (other weights remain $\lambda_{seg}=1.0,\lambda=1.0$). EMA decay is increased to $\alpha=0.999$, and the segmentation confidence threshold is raised to $\tau=0.95$. We use a mini-batch size of 32 (16 labeled, 16 unlabeled) and train for 120,000 iterations with a $321\times 321$ crop size.
ADE20K: We adopt the same loss weights ($\lambda_{seg},\lambda_{bdry},\lambda_{dual},\lambda$), EMA decay ($\alpha$), and segmentation threshold ($\tau$) as used for Pascal VOC. However, specific training parameters differ: the base learning rate is $0.01$, mini-batch size is 16 (8 labeled, 8 unlabeled), training duration is 80,000 iterations, and the input crop size is $513\times 513$.
@@ 4.2.1 Cityscapes
Table 1: Comparison with recent state-of-the-art methods on the Cityscapes dataset. All methods are trained using ResNet-50/101 and DeepLabV3+. For our methods, we averaged over three runs. The evaluation metric is the mIoU (%).
Cityscapes ResNet-50 ResNet-101
@@
CVCC [32] [CVPR’23] 74.9 76.4 77.3 - - -
iMAS [38] [CVPR’23] 74.3 77.4 78.1 - - -
@@
PrevMatch [42] [ArXiv’24] 75.8 77.8 78.8 77.7 78.9 80.1
UCCL [36] [ICASSP’25] 75.8 77.2 78.2 - - -
NRCR [43] [NN’25] - 77.0 77.9 - 78.2 79.5
@@
CW-BASS [14] [ArXiv’25] 75.0 77.2 78.4 - - -
[…]
SAMTH+BoundMatch 76.5 77.9 79.3 77.9 79.0 80.1
@@
Sec. 4.2.1 summarizes the quantitative results on the Cityscapes dataset across various labeled data splits ($\nicefrac{{1}}{{16}}$, $\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{4}}$) for both ResNet-50 and ResNet-101 backbones. Compared with recent state-of-the-art methods, SAMTH+BoundMatch achieves significant performance improvements, particularly in the low-data regime. Notably, incorporating the boundary-aware components allows SAMTH+BoundMatch to consistently outperform the SAMTH baseline, yielding mIoU gains ranging from 1.4% to 2.6% across the tested settings. We attribute this improvement, especially pronounced with limited labeled data, to the additional regularization provided by the boundary multi-task learning objective, as strong regularization techniques are often particularly effective in low-supervision scenarios. The qualitative results presented in Fig. 5 further illustrate that SAMTH+BoundMatch produces more precise object boundaries and improves segmentation accuracy for challenging classes such as “wall”, “pole”, “train”, and “bus”. Overall, these findings establish SAMTH as a strong baseline for urban scene SS-SS, while SAMTH+BoundMatch demonstrates competitive performance against state-of-the-art methods through its effective integration of boundary awareness.
@@ 4.2.2 BDD100K, SYNTHIA, and ADE20K
Table 2: Comparison with our method (SAMTH+BoundMatch) against UniMatch and PrevMatchon the BDD100K, SYNTHIA, and ADE20K datasets. All methods use DeepLabV3+ with ResNet-50 backbone.
@@
PrevMatch [42] - - - - - - 15.4 19.6 24.9
[…]
To assess the generalizability of our approach, we extended the evaluation to additional datasets: the urban driving scene datasets BDD100K and SYNTHIA, as well as the challenging academic benchmark ADE20K, known for its diversity and difficulty in semantic segmentation. As shown in Tab. 2, SAMTH+BoundMatch consistently yields significant improvements over supervised baselines and UniMatch on these datasets. For instance, on BDD100K, our method improves the mean IoU by over 2% for the $\nicefrac{{1}}{{64}}$ split. On the complex ADE20K dataset, our method is competitive against the capable PrevMatch. These results demonstrate that our boundary-aware learning framework effectively leverages unlabeled data across various domains, including challenging datasets with complex scenes.
@@ 4.2.3 Pascal VOC 2012
Table 3: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Classic splits. All methods are trained using ResNet-50/101 and DeepLabV3+.
@@
PrevMatch [42] [ArXiv’24] 73.4 75.4 77.5 77.0 78.5 79.6
UCCL [36] [ICASSP’25] - 74.1 77.1 - - -
@@
CW-BASS [14] [ArXiv’25] 72.8 75.8 76.2 - - -
[…]
UniMatch+BoundMatch 74.2 75.8 76.9 76.0 78.2 78.9
PrevMatch+BoundMatch 74.5 75.8 77.7 77.2 78.7 79.7
Table 4: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Blender splits. All methods are trained using ResNet-50 and DeepLabV3+.
@@
UniMatch [9] [CVPR’23] 75.8 76.9 76.8
CFCG [13] [ICCV’23] 75.6 77.6 78.3
@@
[…]
UniMatch+BoundMatch 76.6 77.9 78.9
PrevMatch+BoundMatch 76.6 77.8 78.0
We further evaluate our method on the Pascal VOC 2012 dataset using both the Classic and Blender splits, acknowledging the challenges posed by this benchmark.
Classic Split: Sec. 4.2.3 presents the performance comparison on the Classic splits for both ResNet-50 and ResNet-101 backbones. It is important to note that Pascal VOC 2012 boundary annotations are known to be noisy, and boundary regions are typically treated as “ignore labels” during training and evaluation. While SAMTH+BoundMatch improves performance relative to earlier CR methods like AugSeg and UniMatch, it does not reach the state-of-the-art results achieved by the most recent approaches, such as PrevMatch and CW-BASS, on this dataset. We hypothesize this performance gap is partly attributable to the weak CR approach of SAMTH, making it less effective for this dataset. To further assess BoundMatch’s boundary mechanism independently of our SAMTH baseline, we integrated it with established CR methods: UniMatch and PrevMatch. When combined with these strong baselines, BoundMatch consistently yields mean IoU gains over the original methods, achieving results that are competitive with, or reach, current state-of-the-art performance (Sec. 4.2.3).
Blended Split: This split contains potentially even more noisy GT labels, particularly around object boundaries. Sec. 4.2.3 shows trends similar to the Classic split. Again, the combinations of BoundMatch with UniMatch or PrevMatch achieve performance competitive with recent SOTA methods. Notably, BoundMatch variants (including SAMTH+BoundMatch and its combinations with UniMatch/PrevMatch) generally outperform CFCG [13], another SS-SS method that utilizes boundary information.
Collectively, these Pascal VOC results underscore the benefit of incorporating explicit boundary modeling in SS-SS frameworks. They demonstrate advantages even when reliant on datasets with imperfect boundary annotations, particularly when BoundMatch’s components enhance already strong CR baselines.

### v2

@@ 3.4 Boundary Consistency Regularized Multi-Task Learning (BCRM)
Figure 3: Overview of the BoundMatch architecture processing labeled samples, illustrating multi-task learning (MTL) of boundary detection alongside segmentation. The Boundary-Semantic Fusion (BSF) module integrates learned boundary cues from the boundary head into the segmentation head’s features, promoting sharper segmentation boundaries. Concurrently, the Spatial Gradient Fusion (SGF) module refines the boundary predictions by fusing them with the spatial gradient of the segmentation mask ($\nabla M$). This refinement improves the quality of boundary supervision for the labeled loss and is crucial for generating cleaner boundary pseudo-labels used in the consistency regularization process for unlabeled data.
@@ 3.4.1 Notation and Boundary Detection Head
We adopt a lightweight boundary detection head inspired by SBCB [11] and CASENet [26]. The architecture leverages hierarchical features from the backbone (e.g. features from Stem, layer1, layer2, and layer4 of a ResNet). Features from the earlier layers are processed to yield single-channel outputs, while the layer4 feature produces an output with $C$ channels, where $C$ is the number of classes. Specifically, the features from these layers are passed through a single $1\times 1$ convolutional layer, respectively, each followed by BN and ReLU activation. These processed features are then bilinearly upsampled to $\nicefrac{{1}}{{2}}$ of the input image size. Following [11], we use an additional $3\times 3$ convolutional layer to reduce artifacts from upsampling. The outputs are then fused using a sliced concatenation and undergo a grouped convolution to form the final boundary prediction map of shape $C\times H\times W$. For more details on the boundary detection head architecture, please refer to Appendix B.1.
Let $Q$ denote the set of boundary predictions (e.g. fused output $q_{\text{fuse}}$ and the layer4 output $q_{\text{last}}$). For the boundary detection task, the ground-truth boundary is denoted by $z$ and the boundary prediction $q$ is a sigmoid-activated probability map in $[0,1]$. We supervise these outputs using a reweighted pixel-wise binary cross-entropy loss:
@@
$\mathcal{L}_{bdry}^{L}=\sum_{q^{S,L}\in Q^{L}}\frac{1}{B^{L}}\sum_{b=1}^{B^{L}}\frac{1}{CHW}\sum_{c,i,j}H_{bdry}\Big{(}z(b,c,i,j),\,q^{S,L}(b,c,i,j)\Big{)}.$ (6)
@@ 3.4.2 MTL for SS-SS with Boundaries
For the labeled set, the boundary detection head is trained alongside the segmentation head. For unlabeled data, we perform consistency regularization analogous to segmentation as shown in Fig. 1. We generate boundary pseudo-labels from the teacher’s weakly augmented predictions $q^{T,w}$, where $q^{T,w}=q_{\text{fuse}}^{T,w}$ represents the teacher’s fused boundary output (as described in Sec. 3.4.1). These are converted to hard pseudo-labels using threshold $\tau_{bdry}$: $\hat{q}^{T}=\mathbf{1}\Big{(}q^{T,w}\geq\tau_{bdry}\Big{)}$. The unsupervised boundary loss then enforces consistency between these pseudo-labels and the student’s strongly augmented predictions:
$\mathcal{L}_{bdry}^{U}\\ =\sum_{q^{S,s}\in Q^{U}}\frac{1}{B^{U}}\sum_{b=1}^{B^{U}}\frac{1}{CHW}\sum_{c,i,j}H_{bdry}\Big{(}\hat{q}^{T}(b,c,i,j),\,q^{S,s}(b,c,i,j)\Big{)}.$ (7)
Note that when SGF is enabled (Sec. 3.5.2), we use the refined boundary prediction $q_{\text{refine}}^{T,w}$ instead of $q_{\text{fuse}}^{T,w}$ to generate higher-quality pseudo-labels, as the refinement process produces cleaner boundaries by incorporating spatial gradient information.
@@ 3.5.1 Boundary-Semantic Fusion (BSF)
Inspired by GSCNN [10], BSF explicitly integrates boundary predictions into the segmentation head. The output from the boundary detection head is concatenated with the features from the Atrous Spatial Pyramid Pooling (ASPP) using the bottleneck layer of the DeepLabV3+ decoder. We emphasize that there are no extra layers introduced by BSF, and we only add extra channels for the bottleneck layer. This explicit incorporation of boundary cues helps localize object boundaries more accurately in the final segmentation.
@@ 3.5.2 Spatial Gradient Fusion (SGF)
Figure 4: Visualization of the output of spatial gradient operator on segmentation prediction.
@@
$[q_{\text{fuse},0},\,q_{\nabla M,0},\,q_{\text{fuse},1},\,q_{\nabla M,1},\,\dots,\,q_{\text{fuse},C-1},\,q_{\nabla M,C-1}],$ (9)
forming a $2C$-channel activation map. This map is then processed by a convolutional layer with $C$-grouped convolutions to yield a refined semantic boundary probability map $q_{\text{refine}}$. This grouped convolution refines the boundary predictions by considering the spatial gradients, effectively sharpening the boundaries between different semantic regions. The refined boundaries are supervised using the same $H_{bdry}$ loss with the labeled set.
Moreover, we introduce a duality loss on the labeled set to enforce consistency between the spatial gradient and the ground-truth boundaries:
@@
We apply this loss only on labeled samples because the ground-truth boundary annotations are reliable; applying it on unlabeled data, which relies on pseudo-labels, introduces instability.
@@ 3.6 BoundMatch
Although demonstrated primarily with our SAMTH baseline throughout the experiments, BoundMatch is designed to be flexible and compatible with various CR methods. For instance, BoundMatch can be integrated into UniMatch [9], a single-model CR method, by incorporating the boundary consistency loss alongside UniMatch’s existing feature-level and strong-view consistency losses. We observed that using only a single model (without a teacher) tends to generate noisy boundary pseudo-labels; therefore, employing an EMA-updated teacher model is crucial for stability and quality.
Similarly, BoundMatch can also be integrated with PrevMatch [44], a state-of-the-art (SOTA) method based on UniMatch, by computing the boundary consistency loss using its robust ensemble pseudo-labels, thereby circumventing potential instability associated with single-iteration predictions.
Furthermore, the architectural simplicity of the BSF and SGF fusion modules facilitates the incorporation of BoundMatch’s boundary-enhancement components into diverse segmentation models, a versatility explored in our ablation studies.
@@ 4.1 Experimental Configuration
Here, we specify the experimental configuration used for evaluating BoundMatch. This includes the datasets utilized (covering urban scenes and academic benchmarks), the model architectures, and key implementation details. We follow the standard SS-SS evaluation protocols for each dataset, ensuring a fair comparison with existing methods following [9].
@@ 4.1.2 Implementation Details
Cityscapes, BDD100K, and SYNTHIA: We use a base learning rate of $0.01$. Loss weights are set to $\lambda_{seg}=1.0$ and $\lambda_{bdry}=1.0$, with an overall unsupervised loss weight $\lambda=1.0$. The EMA decay is $\alpha=0.99$, and the confidence threshold for segmentation pseudo-labels is $\tau=0$. $\tau_{bdry}$ is set to 0.5. Training uses a mini-batch size of 16 (8 labeled, 8 unlabeled) for 80,000 iterations. The input crop size is $801\times 801$ for Cityscapes and $641\times 641$ for BDD100K and SYNTHIA.
Pascal VOC: Hyperparameters are adjusted: base learning rate is $0.001$ and boundary loss weight $\lambda_{bdry}=0.1$ (other weights remain $\lambda_{seg}=1.0,\lambda=1.0$). EMA decay is increased to $\alpha=0.999$, and the segmentation confidence threshold is raised to $\tau=0.95$. $\tau_{bdry}$ is set to 0.5. We use a mini-batch size of 32 (16 labeled, 16 unlabeled) and train for 120,000 iterations with a $321\times 321$ crop size.
ADE20K: We adopt the same loss weights ($\lambda_{seg},\lambda_{bdry},\lambda$), EMA decay ($\alpha$), and thresholds ($\tau$ and $\tau_{bdry}$) as Pascal VOC. However, specific training parameters differ: the base learning rate is $0.01$, mini-batch size is 16 (8 labeled, 8 unlabeled), training duration is 80,000 iterations, and the input crop size is $513\times 513$.
@@ 4.2.1 Cityscapes
Tab. 1 summarizes the quantitative results on the Cityscapes dataset across various labeled data splits ($\nicefrac{{1}}{{16}}$, $\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{4}}$) for both ResNet-50 and ResNet-101 backbones. Compared with recent state-of-the-art methods, SAMTH+BoundMatch achieves significant performance improvements in the low-data regime. Notably, incorporating the boundary-aware components allows SAMTH+BoundMatch to consistently outperform the SAMTH baseline and achieves 1.4% and 2.4% improvements on the difficult $\nicefrac{{1}}{{16}}$ split with ResNet-50 and ResNet-101 respectively. We attribute this improvement to the additional regularization provided by the boundary multi-task learning objective, as strong regularization techniques are often particularly effective in low-supervision scenarios. The qualitative results presented in Fig. 5 further illustrate that SAMTH+BoundMatch produces more precise object boundaries and improves segmentation accuracy for challenging classes such as “wall”, “pole”, “train”, and “bus”. Our proposed simple baseline, SAMTH, also demonstrates strong performance, achieving competitive results against previous SOTA methods. Overall, these findings establish SAMTH as a strong baseline for urban scene SS-SS, while SAMTH+BoundMatch demonstrates competitive performance against state-of-the-art methods through its effective integration of boundary awareness.
As stated previously, BoundMatch can be integrated with other SS-SS methods, and we have shown results for UniMatch [9] and PrevMatch [44] in Tab. 1. For a fair comparison, we reproduced UniMatch and PrevMatch under the same experimental settings as our SAMTH baseline. Incorporating BoundMatch into these methods, we observe consistent performance gains across all splits and backbones. We believe that the performance gains are not as significant as the gains achieved by SAMTH+BoundMatch because UniMatch and PrevMatch already leverage strong consistency regularization techniques which incurs multiple loss functions and may reduce the benefits of the boundary multi-task learning. In future works, we will explore approaches to better balance the regularization signals, but we believe that the results shown here proves the effectiveness of BoundMatch in enhancing SS-SS methods.
Additionally, we reproduced CFCG [13] under our experimental settings to ensure a fair comparison, as the original benchmark partitions may differ from ours. Our results show that SAMTH+BoundMatch outperforms CFCG and the recent CW-BASS [14], highlighting the advantages of our approach. While CFCG and CW-BASS identify boundary regions as difficult areas and adjust loss weighting accordingly (using Laplacian and Sobel operators respectively), they do not aim to improve boundary quality directly. Our method takes a fundamentally different approach: we explicitly learn semantic boundaries through a dedicated auxiliary task and use these learned boundaries not just for identifying difficult regions, but for actively improving both segmentation and boundary predictions through bidirectional fusion modules. This creates a virtuous cycle where better boundaries lead to better segmentation and vice versa.
[…]
ResNet-50 ResNet-101
@@
CVCC [32] [CVPR’23] $74.9$ $76.4$ $77.3$
iMAS [38] [CVPR’23] $74.3$ $77.4$ $78.1$
@@
BoundaryMatch [15] [IS’24] - - - $76.0$ $78.4$ $79.1$
UCCL [36] [ICASSP’25] $75.8$ $77.2$ $78.2$
NRCR [45] [NN’25] - $77.0$ $77.9$ $78.2$ $79.5$
@@
CW-BASS [14] [ArXiv’25] $75.0$ $77.2$ $78.4$
SAMTH (ours) $75.1$ $77.3$ $78.9$ $75.5$ $77.9$ $79.7$
SAMTH + BoundMatch (ours) 76.5 78.1 79.3 77.9 79.0 80.1
+1.4 +0.8 +0.4 +2.4 +1.1 +0.4
UniMatch [9] [CVPR’23] $75.2$ $76.9$ $77.5$ $76.6$ $77.8$ $79.1$
UniMatch + BoundMatch (ours) $76.0$ $77.6$ $78.2$ $77.4$ $78.4$ $79.7$
+0.8 +0.7 +0.7 +0.8 +0.6 +0.6
PrevMatch [44] [ArXiv’24] $75.7$ $77.6$ $78.6$ $77.4$ 78.9 $79.9$
PrevMatch + BoundMatch (ours) 76.3 78.0 $79.1$ 77.8 79.0 80.2
+0.6 +0.4 +0.5 +0.4 +0.1 +0.3
@@ 4.2.2 BDD100K, SYNTHIA, and ADE20K
Table 2: Comparison of SAMTH + BoundMatch with UniMatch on three datasets using DeepLabV3$+$ (ResNet-50).
@@
Ours 52.4 53.9 57.8 69.9 73.0 74.8 15.6 19.6 25.3
To assess the generalizability of our approach, we extended the evaluation to additional datasets: the urban driving scene datasets BDD100K and SYNTHIA, as well as the challenging academic benchmark ADE20K, known for its diversity and difficulty in semantic segmentation. As shown in Tab. 2, SAMTH+BoundMatch consistently yields significant improvements over supervised baselines and UniMatch on these datasets. For instance, on BDD100K, our method improves the mean IoU by over 3% for the $\nicefrac{{1}}{{64}}$ split. These results demonstrate that our boundary-aware learning framework effectively leverages unlabeled data across various domains, including challenging datasets with complex scenes. For qualitative visualizations, please refer to Fig. 13 in the supplementary material, which illustrates the improved segmentation boundaries and overall performance of SAMTH+BoundMatch compared to UniMatch.
@@ 4.2.3 Pascal VOC 2012
We further evaluate our method on the Pascal VOC 2012 dataset using both the Classic and Blender splits, acknowledging the challenges posed by this benchmark. It is important to note that Pascal VOC 2012 boundary annotations are known to be noisy, and boundary regions are typically treated as “ignore labels” during training and evaluation (see Sec. E.2). Therefore, in this benchmark, we hypothesized that the gains from BoundMatch will be limited compared to the previous high-quality datasets. However, experiment results demonstrate that BoundMatch can still provide significant improvements in segmentation performance, especially in the low labeled data regime as shown in this section.
Table 3: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Classic splits. All methods are trained using DeepLabV3$+$ (ResNet-50/101).
@@
BoundaryMatch [15] [IS’24] - - - $75.4$ $77.3$ $79.3$
UCCL [36] [ICASSP’25] - $74.1$ $77.1$
@@
CW-BASS [14] [ArXiv’25] $72.8$ 75.8 $76.2$
SAMTH (ours) $70.7$ $72.1$ $76.1$ $73.2$ $76.4$ $78.5$
SAMTH + BoundMatch (ours) $72.6$ $73.8$ $77.3$ $76.6$ $78.3$ $78.9$
+1.9 +1.7 +1.2 +3.4 +1.9 +0.4
UniMatch [9] [CVPR’23] $71.9$ $72.5$ $76.0$ $75.2$ $77.2$ $78.8$
UniMatch + BoundMatch (ours) 74.2 75.8 $76.9$ $76.0$ $78.2$ $78.9$
+2.3 +3.3 +0.9 +0.8 +1.0 +0.1
PrevMatch [44] [ArXiv’24] $73.4$ 75.4 77.5 $77.0$ $78.5$ $79.6$
PrevMatch + BoundMatch (ours) 74.5 75.8 77.7 77.5 78.7 $79.7$
+1.1 +0.4 +0.2 +0.5 +0.2 +0.1
Table 4: Comparison with recent state-of-the-art methods on the Pascal VOC 2012 dataset using the Blender splits. All methods are trained using DeepLabV3$+$ (ResNet-50).
@@
CFCG [13] [ICCV’23] $75.2$ $76.7$ $77.1$
@@
SAMTH (ours) $73.5$ $75.8$ $76.9$
SAMTH + BoundMatch (ours) $76.3$ $77.2$ $78.1$
+2.8 +1.4 +1.2
UniMatch [9] [CVPR’23] $76.0$ $76.9$ $76.7$
UniMatch + BoundMatch (ours) 76.6 77.9 78.9
+0.6 +1.0 +2.2
PrevMatch [44] [ArXiv’24] $75.8$ $77.0$ $77.7$
PrevMatch + BoundMatch (ours) 76.7 $77.8$ 78.8
+0.9 +0.8 +1.1
Table 5: Comparison with recent state-of-the-art methods using DPT with DINOv2 backbones on the Cityscapes dataset.
Methods DINOv2-S DINOv2-B
$\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$ $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
Supervised $77.2$ $80.2$ $80.8$ $82.7$
UniMatch-V2 [49] [TPAMI’25] $80.6$ $81.9$ $83.6$ $84.3$
SegKC [50] [CoRR’25] $81.2$ $82.4$
SAMTH (ours) $80.8$ $82.0$ $83.2$ $84.1$
SAMTH + BoundMatch (ours) 81.5 82.9 84.0 84.8
+0.7 +0.9 +0.8 +0.7
Classic Split: Tab. 3 presents the performance comparison on the Classic splits for both ResNet-50 and ResNet-101 backbones. While SAMTH+BoundMatch improves performance relative to earlier CR methods like AugSeg and UniMatch, it does not reach the state-of-the-art results achieved by the most recent approaches, such as PrevMatch and CW-BASS, on this dataset. We hypothesize this performance gap is partly attributable to the weak CR approach of SAMTH, making it less effective for this dataset. To further assess BoundMatch’s boundary mechanism independently of our SAMTH baseline, we integrated it with UniMatch and PrevMatch again. When combined with these strong baselines, BoundMatch consistently yields mean IoU gains over the original methods, achieving results that are competitive with, or reach, current state-of-the-art performance (Tab. 3).
Blended Split: This split contains potentially even more noisy GT labels, particularly around object boundaries. Tab. 4 shows trends similar to the Classic split. Again, the combinations of BoundMatch with UniMatch or PrevMatch achieve performance competitive with recent SOTA methods. Furthermore, BoundMatch outperforms CFCG [13] in this split, demonstrating its robustness even in the presence of noisy boundary annotations.
Collectively, these Pascal VOC results underscore the benefit of incorporating explicit boundary modeling in SS-SS frameworks. While BoundMatch does not consistently outperform the best methods on this dataset, it demonstrates its potential to enhance segmentation performance when integrated with strong SS-SS baselines attributed to its simple yet effective boundary-aware learning mechanism.
4.2.4 Recent Benchmarks with Transformer Backbones
While DeepLabV3+ with ResNet backbones has been the standard benchmark for SS-SS methods, recent works have begun exploring vision transformers to leverage their superior representation learning capabilities [49]. Following this direction, we evaluate BoundMatch using DPT [51] with DINOv2 [52] pretrained weights, a foundation model that has demonstrated strong performance across various vision tasks. Adapting BoundMatch to DPT requires minimal modifications, with implementation details provided in Appendix B.2.
As shown in Tab. 5, SAMTH+BoundMatch outperforms current state-of-the-art methods, UniMatch-V2 [49] and SegKC [50], across both DINOv2-S and DINOv2-B backbones on the challenging $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ splits. These results demonstrate that BoundMatch effectively scales to modern foundation models while maintaining its boundary-aware advantages. Qualitative visualizations are provided in Appendix E.3.